# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ I Introduction
However, recent experimental findings have challenged this notion, reporting that noise correlations are stronger during correct than during incorrect perceptual judgments [34, 35, 32, 36, 37, 38], suggesting that noise correlations benefit perceptual abilities. This has been reported even in experiments in which noise correlations are information-limiting, ruling out that this effect may be due to enhanced encoded information. While some of these results can be understood also neglecting neural dynamics and considering only time-averaged responses [39], they have a natural explanation considering the time-dependent nature of neural correlations and neural information encoding and transmission. Temporally-correlated presynaptic activity can elicit stronger responses in postsynaptic neurons with short integration time constants via the non-linear mechanism of coincidence detection [24, 40]. Under such conditions, noise correlations can amplify responses in downstream neural populations and thus increase the amount of sensory information that is transmitted downstream, despite reducing the encoded input information. In particular, stronger information-limiting noise correlations may make different neurons more redundant, which enhances the consistency of stimulus-encoding neural responses across time and neurons. A downstream neuron with short integration time constants may exploit the enhanced consistency in its input to outweigh the information-limiting effects of noise correlations, ultimately increasing the stimulus information that is read out downstream [32, 41].
@@
Understanding the feasibility of such an enhanced-by-coincidence information transmission between pre- and post-synaptic activity requires models of neurons as physical systems that integrate both stimulus encoding and readout mechanisms. A biologically plausible Neural Encoding and Readout Model (NERM) was proposed by Valente et al. (2021) [32]. This model consists of a population of pre-synaptic neurons encoding a binary stimulus, which subsequently drives a single post-synaptic readout neuron. The encoding neurons generate spike trains with a certain strength and time scales of correlation across neurons. These correlations are designed to be information-limiting, that is to decrease the information encoded in the input when the strength of correlation increases. The readout neuron is modeled as a leaky integrate-and-fire (LIF) neuron and integrates afferent currents as a weighted sum of spike trains with a given membrane integration time constant. Numerical solutions of this model reveal that the stimulus discriminability of the readout neuron (as quantified by the signal-to-noise ratio between the difference in stimulus-specific mean firing rates and its variance, SNR) can be, seemingly paradoxically, enhanced by stronger information-limiting correlations for sufficiently short readout time constants and sufficiently low input firing rate. In these conditions correlations amplify the stimulus information transmitted to the readout even when decreasing input information.
@@
This previous study left open important questions about the conditions needed for noise correlations to overcome their information-limiting impact at the readout stage. The study of the model was limited to only $N=2$ input encoding neurons, while information transmission in the brain involves larger neural populations. Moreover, this previous study has not provided an analytical understanding of the interplay between encoding and readout of information in neural populations. To overcome these issues and allow and understanding of these neural operations as a statistical physical system, we develop an analytical framework to characterize the SNR enhancement effect beyond numerical simulations. This frameowrk allows to study encoding in a population with an arbitrary number $N$ of pre-synaptic input neurons with time-dependent correlations, and study how this information is transmitted to a readout LIF neuron. The ability to obtain solutions for arbitrary input population sizes allows us to study the interplay between the spatial dimension of population size (including the large $N$ limit) and the time scales of input correlations and of readout integration. Our approach employs a diffusion approximation, where the readout neuron membrane potential obeys a second-order stochastic differential equation. In this framework, the inter-spike interval (ISI) of the readout neuron corresponds to the first passage time of an inertial stochastic process. We derive analytical expressions for the mean and variance of the readout neuron firing rate, and its temporal correlation structure. Treating this problem within the theory of stochastic processes involves analyzing first-passage times of non-Markovian random walks, which typically requires solving coupled Fokker-Planck equations. By introducing a novel analytical approach —the quenched-noise approximation— we provide tractable expressions for the firing rate mean and variance under the assumption that the noise correlations timescale ${\tau_{\rm c}}$ is much larger than the membrane time constant of the readout neuron ${\tau_{\rm m}}$. The quenched-noise approximation exploits classical expressions for the mean rate of the LIF neuron with white noise input [42]. Our results extend previous expressions for the mean rate of LIF neurons with colored noise input with ${\tau_{\rm m}}\ll{\tau_{\rm c}}$ [17], which were recovered in the limit of weak correlation strength. Moreover, we are able to study analytically the opposite limit ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ using a Poisson approximation. These results provide the necessary conditions for noise correlations to enhance the readout SNR. Notably, our analysis demonstrates that SNR enhancement can occur for a wide range of noise correlation timescales ${\tau_{\rm c}}$, regardless of whether they exceed or fall below the membrane time constant ${\tau_{\rm m}}$.
@@
The article is structured as follows. We first define the NERM (Sec. II). We then (Sec. III) investigate how a readout LIF neuron with or without time-correlated inputs discriminates between binary stimuli $s$ by computing the SNR of its firing activity. In Sec. III.1, we approximate analytically the SNR using the mean and variance of the firing rate of a LIF neuron with input current satisfying a scalar Ornstein-Uhlenbeck equation, within the quenched-noise approximation in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ regime (Secs. III.2, III.4). In Sec. III.3, we explain intuitively why and when noise correlations enhance downstream stimulus information transmission under the quenched-noise approximation. In Sec. IV we describe analytically how noise correlations influence the readout SNR. We finally draw conclusions in Sec. V.
@@ II Definition of the Neural Encoding-Readout Model (NERM)
To study the impact of noise correlations at the encoding (input, pre-synaptic) stage on a post-synaptic readout neuron, we consider the biophysical Neural Encoding-Readout Model (NERM) defined in [32]. The NERM consists of a population of ${N}$ neurons encoding a binary stimulus $s=\pm 1$, and a single readout neuron that integrates the neural activity of the encoding population (see Figs. 1,2,3).
@@
where ${\bm{\eta}}(t)$ is the vector of across-time and across-neuron uncorrelated white noises, ${\bm{\mu}}^{(s)}$ the stimulus-specific mean activity (or tuning curve), $C$ the stationary across-neuron covariance matrix and ${\tau_{\rm c}}$ the correlation timescale (see Table 2 in Appendix A). Thus, the resulting firing rates are correlated across neurons with covariance $C$ and across time with timescale ${\tau_{\rm c}}$, with first and second moments given by $\langle x_{j}^{(s)}(t)\rangle_{\bm{\eta}}=\mu_{j}^{(s)}$ and $\langle(x_{i}^{(s)}(t^{\prime})-\mu_{i}^{(s)})(x^{(s)}_{j}(t^{\prime}+t)-\mu_{j}^{(s)})\rangle_{\bm{\eta}}=C_{ij}e^{-|t|/{\tau_{\rm c}}}$, where the expectation values $_{\bm{\eta}}$ are computed across white noise realizations [43, 44]. We define the matrix of noise correlations, $\bar{C}$, as $\bar{C}_{ij}:=C_{ij}/({v_{\rm V}}_{i}{v_{\rm V}}_{j})^{1/2}$, where ${v_{\rm V}}_{j}:=C_{jj}$ is the stationary variance of the $j$-th neuron rate. The corresponding spike count variance is ${\sigma_{\rm V}}_{j}^{2}=2{\tau_{\rm c}}{v_{\rm V}}^{2}$ (see Appendix A for a discussion about the stationary vs spike count variance terminology). The noise correlation matrix is defined to exhibit equal off-diagonal elements, i.e. $\bar{C}_{ij}={\alpha_{\rm V}}^{2}$ for $i\neq j$, where the fraction of shared noise ${\alpha_{\rm V}}\in[0,1]$ modulates the shared over private noise ratio, so that the limiting cases ${\alpha_{\rm V}}=0,1$ represent completely uncorrelated and correlated encoding firing rates respectively. To ensure that the noise correlations are information-limiting, we set the noise amplitude ${\sigma_{\rm V}}_{i}$ to be proportional to the difference of tuning curves ${\sigma_{\rm V}}_{i}={\nu_{\rm V}}\,{\Delta\mu}_{i}$, where ${\Delta\mu}_{i}:=\mu_{i}^{(+)}-\mu_{i}^{(-)}>0$ with ${\nu_{\rm V}}>0$ (see [32] and equation (8) below).
@@
Examples of the dynamics of the input spike trains at the encoding stage and of the membrane and spiking dynamics of the readout neuron are given in Fig 2 for the case of correlated inputs ($\alpha_{V}=0.9$) and in Fig. 3 for the case when correlations in the input are set to zero ($\alpha_{V}=0$). Note that the input activity of the model with $\alpha_{V}=0.9$ exhibits clearly visible covariations over the 100 ms time-scale set by the Ornstein-Uhlenbeck process (Fig 2) compared to the $\alpha_{V}=0$ case. The output of the readout neuron has higher firing and higher rate separation (and thus higher stimulus-discriminative signal) between the rates elicited by the two stimuli, because correlations in inputs increase the frequency of events with near-coincident firing in the input compared to the uncorrelated-input case (Fig. 3). However, input correlations also increase the variability, or noise, of the readout neuron output rate, which makes it difficult to understand whether input correlations increase or decrease the output information without a careful quantification of both signal and noise as function of the parameters of the model.
To examine how well the stimulus can be discriminated from the readout spike trains, we consider, as in [32], the (square root of the) Signal-to-Noise Ratio (SNR), defined as:
$\text{SNR}=\frac{{r_{\rm R}}^{(+)}-{r_{\rm R}}^{(-)}}{\left[({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2\right]^{1/2}},$ (4)
where ${r_{\rm R}}^{(s)}$ and ${v_{\rm R}}^{(s)}$ are the stimulus-specific average and variance of the readout rate ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$ over the distribution of readout spike trains $\vec{t}$, i.e. $\text{prob}(\vec{t}\,|s)$ (see Table 2). In our numerical analyses, we focus on the SNR in the unit time interval $\text{SNR}\times{T_{\rm c}}^{-1/2}$, since the rate variance decreases as ${T_{\rm c}}^{-1}$ (Eq. (11)) and the SNR increases as ${T_{\rm c}}^{1/2}$.
@@
Figure 2: Neural activity in the Neural Encoding and Readout Model with ${N}=2$ encoding neurons and a strong degree of noise correlation alignment ${\alpha_{\rm V}}=0.9$. (A-B) Encoding firing rates $x^{(s)}_{1,2}(t)$ (continuous curves in A,B) and spike trains sampled from an inhomogeneous Poisson process with rate $x^{(s)}_{1,2}(t)$ are shown in a time window $[0,T=4{\rm s}]$ for stimuli $s=\pm 1$ (orange and violet, respectively). (C-D) LIF membrane potential $V(t)$ (grey lines) and spike times (ticks) for $s=+,-$ (respectively panels C and D). (E) Encoding spike density $(n_{1}(\Delta),n_{2}(\Delta))$ in different time windows of size $\Delta=0.5$ (for a simulation that lasts up to ${T_{\rm c}}=100$). The ellipse axes are oriented along the eigenvectors of $\bar{C}$, and their lengths are proportional to the corresponding eigenvalue square roots. (F) Average (diamonds) and standard deviation (error-bars) of the encoding rate density. While the averages are simply $\sum_{i}\mu_{i}^{(\pm)}$, the standard deviations are their theoretical value, $({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})/w$. The inset shows the theoretical value (continuous lines), along with the numerical estimation of the rate standard deviation, by means of the Jack-Knife method [47], using an increasing sliding window $\Delta_{\rm J}$ (abscissa). (G) Average (circles) and standard deviation (error-bars) of the readout neuron firing rate (${T_{\rm c}}=100$ in simulations). The difference in the averages divided by the (stimulus-averaged) error is the SNR. In this figure, we estimated the standard deviation by means of the Jackknife error (inset ordinates) for increasing sliding window (inset abscissa), and taking the largest value. The simulation parameters are: $N=2$, ${\alpha_{\rm V}}=0.9$, $\sigma_{i}=12$, ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.05$, ${\mu_{\rm I}}^{(-)}=100$, ${\mu_{\rm I}}^{(+)}=120$, $w=0.05$. See [48] and Appendix C for further details.
Figure 3: Neural activity in the neural encoding and readout model, as in Fig. 2, but in absence of noise correlations ${\alpha_{\rm V}}=0$. Note the agreement between the numerical and theoretical rate standard deviation in panel F. Note as well that, even in absence of off-diagonal noise correlations ${\alpha_{\rm V}}=0$, the total firing rate $x^{(s)}(t)$ oscillates around its mean (because the diagonal of the noise correlation matrix is nonzero: ${\sigma_{\rm c}}\neq 0$ even if ${\alpha_{\rm V}}=0$ in Eq. (6c)). We can see that the encoding SNR increases, with respect to the prevalence of noise correlations in Fig. 2 (since $\mu^{(+)}-\mu^{-}$ is constant, while ${\sigma_{\rm c}}$ has decreased). The rest of the parameters are: $N=2$, ${\alpha_{\rm V}}=0$, $\sigma_{i}=12$, ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.05$, ${\mu_{\rm I}}^{(-)}=100$, ${\mu_{\rm I}}^{(+)}=120$, $w=0.05$.

### v2

@@ I Introduction
However, recent experimental findings, reviewed in the next section and sketched in Fig.1, have challenged this notion, reporting that noise correlations are stronger during correct than during incorrect perceptual judgments [34, 35, 32, 36, 37, 38], suggesting that noise correlations benefit perceptual abilities. This has been reported even in experiments in which noise correlations are information-limiting, ruling out that this effect may be due to enhanced encoded information. While some of these results can be understood also neglecting neural dynamics and considering only time-averaged responses [39], they have a natural explanation considering the time-dependent nature of neural correlations and neural information encoding and transmission. Temporally-correlated presynaptic activity can elicit stronger responses in postsynaptic neurons with short integration time constants via the non-linear mechanism of coincidence detection [24, 40]. Under such conditions, noise correlations can amplify responses in downstream neural populations and thus increase the amount of sensory information that is transmitted downstream, despite reducing the encoded input information. In particular, stronger information-limiting noise correlations may make different neurons more redundant, which enhances the consistency of stimulus-encoding neural responses across time and neurons. A downstream neuron with short integration time constants may exploit the enhanced consistency in its input to outweigh the information-limiting effects of noise correlations, ultimately increasing the stimulus information that is read out downstream [32, 41].
@@
Understanding the feasibility of such an enhanced-by-coincidence information transmission between pre- and post-synaptic activity requires models of neurons as physical systems that integrate both stimulus encoding and readout mechanisms. To understand how correlated input at the pre-synaptic stage enhance information transmission to the post-synaptic stage, we propose a biologically plausible Neural Encoding and Readout Model (NERM)[32]. The model consists of a population of pre-synaptic neurons encoding a binary stimulus, which subsequently drives a single post-synaptic readout neuron. The encoding neurons generate spike trains with a certain strength and time scales of correlation across neurons. These correlations are designed to be information-limiting, that is to decrease the information encoded in the input when the strength of correlation increases. The readout neuron is modeled as a leaky integrate-and-fire (LIF) neuron and integrates afferent currents as a weighted sum of spike trains with a given membrane integration time constant. Numerical solutions of this model reveal that the stimulus discriminability of the readout neuron (as quantified by the signal-to-noise ratio between the difference in stimulus-specific mean firing rates and its variance, SNR) can be, seemingly paradoxically, enhanced by stronger information-limiting correlations for sufficiently short readout time constants and sufficiently low input firing rate. In these conditions correlations amplify the stimulus information transmitted to the readout even when decreasing input information.
@@
Our previous study left open important questions about the conditions needed for noise correlations to overcome their information-limiting impact at the readout stage. The study of the model was limited to only $N=2$ input encoding neurons, while information transmission in the brain involves larger neural populations. Moreover, we did not provide an analytical understanding of the interplay between encoding and readout of information in neural populations. To overcome these issues and allow an understanding of these neural operations as a statistical physical system, here we develop an analytical framework to characterize the SNR enhancement effect beyond numerical simulations. This framework allows to study encoding in a population with an arbitrary number $N$ of pre-synaptic input neurons with time-dependent correlations, and study how this information is transmitted to a readout LIF neuron. The ability to obtain solutions for arbitrary input population sizes allows us to study the interplay between the spatial dimension of population size (including the large $N$ limit) and the time scales of input correlations and of readout integration. Our approach employs a diffusion approximation, where the readout neuron membrane potential obeys a second-order stochastic differential equation. In this framework, the inter-spike interval (ISI) of the readout neuron corresponds to the first passage time of an inertial stochastic process. We derive analytical expressions for the mean and variance of the readout neuron firing rate, and its temporal correlation structure. Treating this problem within the theory of stochastic processes involves analyzing first-passage times of non-Markovian random walks, which typically requires solving coupled Fokker-Planck equations. By introducing a novel analytical approach —the quenched-noise approximation— we provide tractable expressions for the firing rate mean and variance under the assumption that the noise correlations timescale ${\tau_{\rm c}}$ is much larger than the membrane time constant of the readout neuron ${\tau_{\rm m}}$. The quenched-noise approximation exploits classical expressions for the mean rate of the LIF neuron with white noise input [42]. Our results extend previous expressions for the mean rate of LIF neurons with colored noise input with ${\tau_{\rm m}}\ll{\tau_{\rm c}}$ [17], which were recovered in the limit of weak correlation strength. Moreover, we are able to study analytically the opposite limit ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ using a Poisson approximation. These results provide the necessary conditions for noise correlations to enhance the readout SNR. Notably, our analysis demonstrates that SNR enhancement can occur for a wide range of noise correlation timescales ${\tau_{\rm c}}$, regardless of whether they exceed or fall below the membrane time constant ${\tau_{\rm m}}$.
@@
The article is structured as follows. We first review experimental work and define the resulting NERM (Sec. II). We then (Sec. III) investigate how a readout LIF neuron with or without time-correlated inputs discriminates between binary stimuli $s$ by computing the SNR of its firing activity. In Sec. III.1, we approximate analytically the SNR using the mean and variance of the firing rate of a LIF neuron with input current satisfying a scalar Ornstein-Uhlenbeck equation, within the quenched-noise approximation in the ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ regime (Secs. III.2, III.4). In Sec. III.3, we explain intuitively why and when noise correlations enhance downstream stimulus information transmission under the quenched-noise approximation. In Sec. IV we describe analytically how noise correlations influence the readout SNR. We finally draw conclusions in Sec. V.
@@ II Definition of the Neural Encoding-Readout Model (NERM)
Before defining our work, for self-consistence of this study we briefly summarize (and schematize in a summary cartoon in Fig.1) the biological experimental and modeling work that leads to it. Experimental work has shown that noise correlations in local brain circuits are predominantly information-limiting [29, 26, 30, 31, 32], as simultaneously-recorded activity (which contains correlations between the activity of different neurons) encodes less information that when this activity is randomly shuffled across trials to form pseudo-population shuffled responses that have the same single-neuron probabilities as the real data but no correlations across neurons (Fig.1A). However, the same data also shows stronger across-neuron correlations in trials in which animals make behaviorally correct judgment about the stimulus than in error trials in which animals make incorrect judgments [34, 35, 32, 36, 37, 38]. This suggests that these correlations may enhance the readout of correlated activity and aid task performance, even when hindering encoded information (Fig.1B). Previous work reconciled these observations using a biophysical Neural Encoding-Readout Model (NERM), where the readout integrated correlated activity from two input units with a short membrane time constant [32] (see Fig.1C). To study the impact of noise correlations at the encoding (input, pre-synaptic) stage on a post-synaptic readout neuron, we extend the NERM to a population of ${N}$ neurons encoding a binary stimulus $s=\pm 1$, and a single readout neuron that integrates the neural activity of the encoding population (see Figs. 2,3,4).
@@
where ${\bm{\eta}}(t)$ is the vector of across-time and across-neuron uncorrelated white noises, ${\bm{\mu}}^{(s)}$ the stimulus-specific mean activity (or tuning curve), $C$ the stationary across-neuron covariance matrix and ${\tau_{\rm c}}$ the correlation timescale (see Table 2 in Appendix A). Thus, the resulting firing rates are correlated across neurons with covariance $C$ and across time with timescale ${\tau_{\rm c}}$, with first and second moments given by $\langle x_{j}^{(s)}(t)\rangle_{\bm{\eta}}=\mu_{j}^{(s)}$ and $\langle(x_{i}^{(s)}(t^{\prime})-\mu_{i}^{(s)})(x^{(s)}_{j}(t^{\prime}+t)-\mu_{j}^{(s)})\rangle_{\bm{\eta}}=C_{ij}e^{-|t|/{\tau_{\rm c}}}$, where the expectation values $_{\bm{\eta}}$ are computed across white noise realizations [43, 44]. We define the matrix of noise correlations, $\bar{C}$, as $\bar{C}_{ij}:=C_{ij}/({v_{\rm V}}_{i}{v_{\rm V}}_{j})^{1/2}$, where ${v_{\rm V}}_{j}:=C_{jj}$ is the stationary variance of the $j$-th neuron rate. The corresponding spike count variance is ${\sigma_{\rm V}}_{j}^{2}=2{\tau_{\rm c}}{v_{\rm V}}_{j}$ (see Appendix A for a discussion about the stationary vs spike count variance terminology). The noise correlation matrix is defined to exhibit equal off-diagonal elements, i.e. $\bar{C}_{ij}={\alpha_{\rm V}}^{2}$ for $i\neq j$, where the fraction of shared noise ${\alpha_{\rm V}}\in[0,1]$ modulates the shared over private noise ratio, so that the limiting cases ${\alpha_{\rm V}}=0,1$ represent completely uncorrelated and correlated encoding firing rates respectively. To ensure that the noise correlations are information-limiting, we set the noise amplitude ${\sigma_{\rm V}}_{i}$ to be proportional to the difference of tuning curves ${\sigma_{\rm V}}_{i}={\nu_{\rm V}}\,{\Delta\mu}_{i}$, where ${\Delta\mu}_{i}:=\mu_{i}^{(+)}-\mu_{i}^{(-)}>0$ with ${\nu_{\rm V}}>0$ (see [32] and equation (8) below).
@@
Examples of the dynamics of the input spike trains at the encoding stage and of the membrane and spiking dynamics of the readout neuron are given in Fig 3 for the case of correlated inputs ($\alpha_{V}=0.9$) and in Fig. 4 for the case when correlations in the input are set to zero ($\alpha_{V}=0$). Note that the input activity of the model with $\alpha_{V}=0.9$ exhibits clearly visible covariances over the 100 ms time-scale set by the Ornstein-Uhlenbeck process (Fig 3-A,B) compared to the $\alpha_{V}=0$ case. The output of the readout neuron exhibits higher rate separation $\Delta{r_{\rm R}}$ (Fig. 3-G), and higher stimulus-discriminative signal between the rates elicited by the two stimuli, because correlations in inputs increase the frequency of events with near-coincident firing in the input compared to the uncorrelated-input case (Fig. 4). However, input correlations also increase the variability, or noise, of the readout neuron output rate, which makes it difficult to understand whether input correlations increase or decrease the output information without a careful quantification of both signal and noise as function of the parameters of the model.
To examine how well the stimulus can be discriminated from the readout spike trains, we consider the (square root of the) Signal-to-Noise Ratio (SNR) [32], defined as (mind that $\Delta{\mu_{\rm I}}>0$ and hence ${r_{\rm R}^{(-)}}<{r_{\rm R}^{(+)}}$):
$\text{SNR}({T_{\rm c}})=\frac{{r_{\rm R}}^{(+)}-{r_{\rm R}}^{(-)}}{\left[({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2\right]^{1/2}},$ (4)
where ${r_{\rm R}}^{(s)}$ and ${v_{\rm R}}^{(s)}$ are the stimulus-specific average and variance of the readout rate ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$ over the distribution of readout spike trains $\vec{t}$, i.e. $\text{prob}(\vec{t}\,|s)$ (see Table 2). From now on, we will define the SNR in the unit time interval $\text{SNR}:=\text{SNR}({T_{\rm c}})\times{T_{\rm c}}^{-1/2}$, since the rate variance decreases as ${T_{\rm c}}^{-1}$ (Eq. (11)) and hence $\text{SNR}({T_{\rm c}})$ increases as ${T_{\rm c}}^{1/2}$.
Although Figs. 3,4 have an illustrative scope, the estimation of the NERM SNR with parameters used in these figures, shows (see Sec. D.2 in Appendix D) that the introduction of noise correlations in Fig. 3 significantly enhances the SNR with respect to the case with no noise correlations, and the same value of the rest of the parameters (Fig. 4).
Figure 1: Summary of recent neuro-biological results on the role of neural correlations in information encoding and readout. A. Left: Single-trial neural activity of two neurons in the case of large and small noise correlations. Right: Input information decreases with the strength of noise correlations, while the consistency of input information (fraction of trials where the two neurons signal the same stimulus) increases with the noise correlation strength. B. Schematics of the enhanced-by-consistency readout model in [32] in which the probability that the choice equals the stimulus is higher for trials with consistent input information. This readout accounts for experimental observations [32] of enhanced behavioral performance in presence of correlations and higher noise correlation strength in correct than in error trials. C. Biophysical information-transmission model that implements an enhanced-by-consistency readout. When the readout unit has a sufficiently short membrane time-constant correlated input activity can drive the readout to spike. Under this condition the readout transmits more stimulus information when input noise correlations are strong, so that increasing noise correlations increases the amount of transmitted information, even when decreasing the amount of input stimulus information.
@@
Figure 3: Neural activity in the Neural Encoding and Readout Model with ${N}=6$ encoding neurons and a strong degree of noise correlation alignment ${\alpha_{\rm V}}=0.9$. (A-B) Encoding firing rates $x^{(s)}_{1,2}(t)$ (continuous curves in A,B) and spike trains sampled from an inhomogeneous Poisson process with rate $x^{(s)}_{1,2}(t)$ are shown in a time window $[0,T=4{\rm s}]$ for stimuli $s=\pm 1$ (orange and violet, respectively). (C-D) LIF membrane potential $V(t)$ (lines) and spike times (ticks) for $s=-,+$ (respectively panels C and D). (E) Scatter plot of encoding spike densities $(\rho_{1},\rho_{2})$, with $\rho_{j}=n_{j}(\Delta)/\Delta$. Each point corresponds to a different time window of size $\Delta_{\rm e}=0.05$, for a simulation that lasts up to ${T_{\rm c}}=200$. The ellipse’s largest axis is oriented along the $(1,1)$ axis, and the larger and lower axes of the ellipse are set to $\nu$ and $(1-{\alpha_{\rm V}}^{2})^{1/2}\nu$ respectively, where $\nu={\sigma_{\rm V}}+\bar{\mu}^{1/2}$ is the theoretical noise amplitude in the diffusion approximation. (F) Average (diamonds) and standard deviation (error-bars) of the encoding rate density. While the averages are simply $w\sum_{i}\mu_{i}^{(\pm)}$, the standard deviations are their theoretical value, $[({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})/w]^{1/2}$. The inset shows the theoretical value (continuous lines) of the rate standard deviation, along with its numerical estimation by means of the Jack-Knife method [47], using an increasing sliding window $\Delta_{\rm J}$ (in abscissa). (G) Box-plots of the histogram of readout neuron firing rates, across ${\rm int}({T_{\rm c}}/\Delta_{\rm r})$ different time intervals of length $\Delta_{\rm r}=0.5$. The difference in the averages divided by the (stimulus-averaged) error is the SNR. The rate standard deviation is estimated by means of the Jackknife error (inset ordinates) for increasing sliding window (inset abscissa). See details on this simulation in Sec. D.2 in Appendix D. The simulation parameters are: ${N}=6$, ${\alpha_{\rm V}}=0.9$, ${\sigma_{\rm V}}=8.9442$, ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.05$, ${\mu_{\rm I}}^{(-)}=80$, ${\mu_{\rm I}}^{(+)}=90$, $w=0.1$. See [48] and Appendix D.2 for further details.
Figure 4: Neural activity in the neural encoding and readout model, as in Fig. 3, but in absence of noise correlations ${\alpha_{\rm V}}=0$. Note that, even in absence of off-diagonal noise correlations ${\alpha_{\rm V}}=0$, the total firing rate $x^{(s)}(t)$ oscillates around its mean (because the diagonal of the noise correlation matrix is nonzero: ${\sigma_{\rm c}}\neq 0$ even if ${\alpha_{\rm V}}=0$ in Eq. (6c)). We can see that the encoding SNR increases, with respect to the prevalence of noise correlations in Fig. 3 (since $\mu^{(+)}-\mu^{-}$ is constant, while ${\sigma_{\rm c}}$ has decreased). The rest of the parameters are: $N=6$, ${\alpha_{\rm V}}=0$, ${\sigma_{\rm V}}=8.9442$, ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.05$, ${\mu_{\rm I}}^{(-)}=80$, ${\mu_{\rm I}}^{(+)}=90$, $w=0.1$.